# LibMoE: A Library for Comprehensive Research Mixture of Experts in Large Language Models

### v2

@@ 3.3 Pre-training LLM: Decoder-only Switch Transformer
A key design of our language modeling experiment is to establish a standardized yet flexible setup that enables fair comparison across different MoE algorithms while maintaining computational efficiency. We adopt a decoder-only Transformer architecture in the style of Switch Transformer (Fedus et al., 2022), where the feedforward sublayer is replaced by sparsely activated Mixture-of-Experts (MoE) modules. This design supports scaling from small (158M) to large (679M) models and provides a unified backbone for evaluating MoE variants. To ensure consistency, we keep the architecture and optimization pipeline fixed across experiments, varying only the routing strategies.
Training pipeline. Our experimental setup builds upon the framework of Csordás et al. (2023), with adjustments tailored to our model configurations and minor refinements to the baseline hyperparameters. The codebase supports all major forms of parallelism, including data, tensor, pipeline, and expert parallelism, ensuring scalability across model sizes. MoE variants are implemented by modularly replacing the MoE component, while keeping other training components identical. This design ensures consistency and reproducibility across experiments, allowing us to isolate the impact of routing. Full specifications of model scales, datasets, and hyperparameters are provided in Appendix A.2 and Appendix B.2.
@@ 3.5 Design Principles
LibMoE is designed with a strong emphasis on modularity, extensibility, and practical usability to accelerate research on Sparse Mixture-of-Experts (SMoE) models. At its core, LibMoE consists of three major modules. The MoE module provides a unified abstraction for routing strategies, supporting customizable designs such as gating mechanisms, load balancing losses, and expert selection logic. The training module handles optimization workflows across both LLM and VLM tasks, allowing researchers to integrate pretrained backbones, apply sparse upcycling or pretraining from scratch, and utilize distributed strategies such as data parallelism and model sharding. Finally, the evaluation module enables seamless plug-and-play evaluation, supporting a wide range of benchmarks and task formats through standardized APIs.
@@ 4.1 SMoE Algorithms
To demonstrate that LibMoE is well-suited for SMoE research, we implement and benchmark seven state-of-the-art SMoE algorithms. First, we consider methods that modify or improve the router network. These include the original SMoE (Fedus et al., 2022), which uses a softmax router; the sigmoid-gated SMoE (Csordás et al., 2023), which has been shown to achieve faster convergence; and XMoE (Chi et al., 2022), which mitigates representation collapse and enhances routing capacity through input dimension reduction and cosine-normalized routing. Second, we incorporate recent designs that utilize shared expert mechanisms, such as DeepSeek-V2 (Liu et al., 2024a) with softmax routing and DeepSeek-V3 (Liu et al., 2024c) with a sigmoid router. Inspired by these, we further propose SharedE-V2 and SharedE-V3, which follow the shared expert design principles of DeepSeek-V2 and DeepSeek-V3, respectively. Lastly, we include MoE++(Jin et al., 2024), which augments standard SMoE with zero-computation experts and pathway-aware routing to improve efficiency and throughput, and TC-MoE(Yan et al., 2025), which introduces a ternary expert choice space and reward-based routing to reduce redundancy while enhancing model expressiveness. These diverse algorithms allow us to evaluate LibMoE’s compatibility with a wide range of routing strategies and architectural innovations.
4.2 Vision-Language Modeling
@@
MoE Method AI2D Text VQA GQA MM Bench Hallusion Bench Math Vista MMMU MMStar Pope MME MME RW OCR Bench $\uparrow AVG$ Rank $\downarrow$ LLAVA + OneVision / 1M2 samples SMoE 69.56 43.93 61.51 71.31 46.90 37.90 41.56 41.23 86.28 63.33 27.83 37.50 52.40 5.96 XMoE 69.72 43.93 61.52 72.25 47.42 38.50 42.11 43.99 86.61 63.81 29.18 39.40 53.20 3.33 $\sigma$-MoE 69.79 44.69 61.70 71.74 47.00 38.40 43.11 42.08 86.69 63.80 29.70 38.40 53.09 3.33 SharedE-V2 70.56 45.04 61.34 71.13 47.11 39.50 42.78 42.73 86.53 63.93 29.55 38.70 53.24 3.33 SharedE-V3 71.92 44.59 61.93 72.59 46.37 38.20 42.33 43.30 86.78 64.61 29.29 39.40 53.44 2.46 TC-MOE 70.08 43.75 61.89 71.05 45.74 38.10 41.89 43.64 86.76 63.05 31.84 38.30 53.01 4.50 MOE++ 70.13 43.37 61.52 71.39 46.16 38.60 40.78 43.24 86.60 63.26 28.19 37.50 52.56 5.08 LLAVA / 665K samples SMoE 65.52 41.51 61.62 72.25 41.75 29.50 41.67 42.24 87.13 61.05 32.15 31.70 50.67 3.42 XMoE 65.84 41.96 61.61 72.16 41.85 31.60 41.67 40.32 86.64 60.95 32.73 33.20 50.88 3.08 $\sigma$-MoE 65.52 42.08 61.74 71.22 40.80 30.30 41.22 41.99 86.64 61.44 32.31 32.50 50.65 3.71 SharedE-V2 64.77 41.96 61.27 71.74 41.85 30.90 43.33 41.56 86.82 60.52 32.88 31.40 50.75 3.67 SharedE-V3 65.58 42.06 61.26 72.42 41.43 30.60 42.44 41.75 86.81 60.93 31.47 32.60 50.86 3.33 TC-MOE 65.50 40.70 61.19 71.22 42.06 29.30 41.22 41.10 86.53 60.30 31.89 33.00 50.34 5.42 MOE++ 65.03 41.44 60.61 71.74 42.69 30.30 43.00 40.32 86.58 60.11 31.32 31.20 50.36 5.38
At Table 1, we report the performances of the SMoE algorithms evaluated in the vision-language modeling (VLM) task, implemented using our LibMoE framework. We observe that no single method consistently outperforms the others across all benchmarks, with the standard deviation approximately 0.16. Although methods such as XMoE, SharedE-V2 and SharedE-V3 may perform slightly better in certain settings, the improvements over the second-best methods are marginal. We argue that such differences may arise from randomness in initialization and stochastic expert routing across SMoE layers, and are likely to diminish as the training data scale increases. Additionally, we extend the experiments to the pre-training setting, as shown in Table 2. In this case, the standard deviation increases to 0.28, partly due to some experts receiving insufficient training data. However, this variation is still not large enough to indicate a truly significant improvement by any single algorithm. Overall, by using LibMoE to compare state-of-the-art SMoE algorithms under a fair and large-scale setting, our findings align with the common consensus that the original SMoE strategy remains the most attractive option due to its simplicity and scalability.
Table 2: Performance comparisons of different Sparse Mixture of Experts (SMoE) models on pre-trained LLM with small-scale model with 158M parameters and large-scale model with 679M parameters. PPL indicates the perplexity score.
@@
MoE Method PPL $\downarrow$ LAM BADA BLiMP CBT Hella Swag PIQA ARC- Easy RACE SIQA Common SenseQA AVG $\uparrow$ AVG Rank $\downarrow$
Dense (36M) 17.04 19.74% 73.48% 81.03% 27.56% 55.82% 32.09% 29.36% 35.72% 25.14% 42.22% 6.78
SMoE 13.63 25.27% 77.71% 84.18% 29.43% 57.94% 32.68% 30.11% 35.62% 24.65% 44.18% 4.56
XMoE 13.98 24.57% 76.53% 84.12% 29.34% 58.27% 32.26% 29.69% 35.47% 24.49% 43.86% 4.44
$\sigma$-MoE 13.61 25.43% 77.38% 84.23% 29.13% 58.92% 32.73% 31.05% 34.90% 24.90% 44.30% 6.56
SharedE-V2 13.49 25.29% 77.37% 84.33% 29.38% 60.17% 33.83% 31.02% 35.57% 24.98% 44.66% 3.00
SharedE-V3 13.42 25.49% 77.20% 84.40% 29.38% 59.14% 32.52% 30.60% 35.57% 25.47% 44.42% 3.44
TC-MoE 13.51 25.60% 76.91% 84.68% 29.27% 59.03% 33.02% 30.63% 36.03% 26.37% 44.62% 2.89
Small Model (158M) MoE++ 13.54 25.45% 77.23% 84.83% 29.28% 58.49% 33.49% 30.11% 35.62% 24.49% 44.33% 4.00
Dense (131M) 11.51 31.00% 77.43% 87.45% 31.90% 61.37% 35.14% 31.58% 36.90% 27.03% 46.64% 7.89
SMoE 9.51 37.13% 80.47% 89.83% 37.49% 64.36% 38.22% 33.03% 37.41% 26.54% 49.39% 5.22
XMoE 9.66 35.25% 80.38% 89.35% 37.19% 64.20% 38.99% 32.95% 37.77% 28.34% 49.38% 3.67
$\sigma$-MoE 9.46 37.56% 81.08% 89.57% 37.52% 64.91% 39.15% 32.68% 37.67% 28.50% 49.85% 5.56
SharedE-V2 9.52 37.11% 80.98% 89.93% 37.14% 64.36% 38.06% 33.17% 36.95% 27.35% 49.45% 5.11
SharedE-V3 9.49 36.88% 81.28% 89.65% 37.32% 65.72% 38.86% 33.12% 38.59% 28.09% 49.95% 3.56
TC-MoE 9.38 37.87% 81.21% 90.19% 37.95% 64.47% 39.28% 33.77% 37.92% 27.85% 50.06% 2.44
Large Model (679M) MoE++ 9.38 38.80% 80.88% 89.77% 37.70% 64.64% 39.37% 34.02% 37.97% 28.34% 50.16% 2.33
Figure 2: Benchmark curves during training in between SMoE and Dense model on pre-training LLM.
4.3 Language Modeling
Table 2 presents the language modeling performance of two model sizes evaluated across multiple methods, implemented using our LibMoE framework. All MoE-based methods significantly outperform the dense counterparts with the same pretraining FLOPs on almost every task, demonstrating the effectiveness of our framework in developing Mixture-of-Experts methods. To further highlight the differences when training between SMoE and Dense models, we illustrate the training curves of two models in Figure 2. We observe that the performance gap between SMoE and Dense further widens through training steps, and SMoE can even match or even outperform dense models trained with $\geq 2\times$ FLOPs.
Our comparative analysis of Mixture-of-Experts (MoE) methods reveals a distinct correlation between performance and the chronological development of these models. Notably, recent approaches such as DeepSeek-V3 (Liu et al., 2024b), MoE++ (Jin et al., 2024), and TC-MoE (Yan et al., 2025) consistently surpass earlier methods, underscoring the progressive advancements achieved by each successive iteration. Models like XMoE (Chi et al., 2022), $\sigma$-MoE (Csordás et al., 2023), and DeepSeek-V2 (Liu et al., 2024b) demonstrate performance that is slightly better or comparable to the Vanilla SMoE (Shazeer et al., 2017b). These results validate the effectiveness of our evaluation framework in benchmarking the evolving landscape of MoE architectures.
@@ 5 Analysis
Beyond the empirical performance, we conduct a comprehensive analysis to uncover how different MoE architectures route, specialize, and balance experts throughout training. Specifically, we aim to disentangle the dynamics of expert selection across algorithms, the degree to which expert capacity is utilized or underexploited, and how factors like initialization, architecture, and task domain shape these behaviors. Our study spans both pretraining from scratch (158M LLM on 6.85B tokens) and sparse upcycling (VLM fine-tuned on LLaVA-665K), offering a unified view of routing behavior across regimes.
@@
This experiment investigates how the amount of training data impacts the expert selection dynamics across various SMoE algorithms. Specifically, we measure the router change rate, defined as the proportion of expert activation decisions that differ between consecutive training checkpoints. A lower change rate reflects more stable routing behavior over time, indicating improved expert specialization and convergence. Figure 3 presents the router change rate throughout training for both vision-language models (MME, MathVista; sparse upcycling) and language models (Validation, HellaSwag; from-scratch pretraining). Across all settings, we observe that SMoE algorithms consistently exhibit decreasing router change rates as training progresses. Among the evaluated methods, XMoE demonstrates the most consistent convergence across both domains in the late stages. We attribute this to its architectural use of cosine-normalized routing, which promotes early expert specialization.
@@
Furthermore, we observe a divergent trend in the behavior of shared expert architectures namely, SharedE-V2 and SharedE-V3 across different tasks. In vision language models (VLMs), both variants exhibit significantly higher routing stability compared to other methods. However, this advantage disappears during language model pretraining. We hypothesize that this divergence arises from the distinct initialization schemes and functional roles of shared experts. In the VLM sparse upcycling setup, SMoE layers are initialized from a pretrained dense model, where shared experts already encode general purpose visual linguistic representations. These pretrained experts act as stable anchors, enabling the router to focus its capacity on distributing tokens among non-shared experts, thereby reducing volatility in expert selection. In contrast, language model pretraining starts from random initialization. Here, both the router and shared experts must co-adapt from scratch. As the shared experts’ representations evolve over time, their influence on the final output changes accordingly. This compels the router to continuously revise its decisions not just for load balancing but also to adapt to the shifting behavior of shared experts in order to minimize the task loss. The result is a tight coupling between unstable shared experts and routing dynamics, which leads to higher expert change rates and delayed convergence.
@@
Taken together, these findings underscore a central insight enabled by LibMoE: the success of sparse MoE architectures depends on distinct factors at different training stages. In language model pretraining, where all components are initialized from scratch, the router and experts must rapidly co-adapt—placing a premium on early-stage stability. In contrast, VLM sparse upcycling benefits from strong shared experts pretrained on dense checkpoints, where routing dynamics evolve more gradually and generalization becomes the priority. As such, designing effective sparse MoE methods requires simultaneously addressing both phases: early adaptation in randomly initialized pretraining and long-term robustness in upcycling regimes.
@@
Figure 5: Performance of SMoE variant when changing top-1 expert to top-$(K+1)$ in vision-language modeling and language modeling task.
To assess whether SMoE algorithms fully exploit their expert capacity, we conduct a diagnostic perturbation experiment across seven architectures. Specifically, we intentionally degrade the routing decision by replacing the top-1 expert (the one with highest affinity score) with the top-(K+1)(K{+}1) expert. This allows us to measure each model’s sensitivity to suboptimal expert assignments and reveal potential routing inefficiencies. Figure 5 reports the cumulative accuracy drop across benchmarks under two settings: replacing only the top-1 expert (DropTop1) and replacing both the top-1 and top-2 experts (DropTop1&\&2). As expected, most SMoE variants exhibit performance degradation, confirming that precise expert selection is critical to maintaining model quality. Overall, in the language modeling (pretraining) setting, all methods suffer large and relatively uniform performance drops, reflecting a consistent reliance on precise expert routing across architectures. In contrast, vision-language models exhibit greater divergence. Notably, SMoE shows a counterintuitive +2.3% gain under the DropTop1 perturbation suggesting that its original routing configuration may have failed to effectively utilize its expert capacity. This exposes a potential inefficiency in the router, where suboptimal expert choices actually yield better outcomes. Conversely, TC-MoE suffers the most severe degradation, highlighting its strong dependence on highly optimized, task-aligned routing decisions.
In summary, this experiment underscores the critical role of routing quality in relation to architectural design across SMoE variants. Architectures such as TC-MoE and MoE++ demonstrate that when expert utilization is effective, performance degrades more sharply under routing perturbations—highlighting the strong coupling between routing robustness and representational efficiency. However, a key insight is that robust routing alone does not guarantee the highest absolute performance. Each SMoE architecture has an inherent capacity ceiling, and some models may approach this limit even with less dynamic or suboptimal routing. In other words, attainable performance is shaped not only by routing precision but also by expert capacity and architectural trade-offs. As such, routing behavior should be interpreted in the broader context of each model’s design constraints and representational limits.
@@
In Figure 6, we examine the behavior of the expert selection mechanism by measuring how frequently each expert is activated across subtasks in the MME benchmark, using the Expert Allocation Entropy (EAE) metric D.1. This analysis enables us to assess the degree of specialization exhibited by individual experts across different domains.
@@
Overall, these findings reveal a spectrum of expert allocation behaviors—from highly specialized but potentially imbalanced (e.g., TC-MoE), to adaptive and task-aware (e.g., XMoE), to balanced yet under-specialized (e.g., SMoE family) highlighting the importance of routing strategies that balance specialization with generalization.
@@
In Figure 7, we report the normalized entropy of expert-weight allocation (EWA), calculated as the Shannon entropy of the softmax router weights over selected experts, and normalized with the maximum entropy of a vector with $K$ elements and averaged over samples within each task group OCR, Coarse-Grained, and Reasoning to reveal trends in how confidently or evenly each routing strategy distributes weight across experts, which EWA larger than weights more uniform. Notably, XMoE exhibits the most desirable profile across all three axes of evaluation. First, it achieves the highest EWA, indicating that the router distributes weight relatively evenly among the selected experts. Second, the entropy remains remarkably consistent across all task groups, with nearly identical scores for Coarse-Grained, Fine-Grained, OCR, and Reasoning. This reflects a routing strategy balance weights to experts reduce appear some expert nominate and each expert contribution to output is nearly equal. At the other extreme, TC-MoE records the lowest EWA across all tasks, with minimal inter-task variation. This indicates that the router consistently concentrates weight on a small subset of experts, regardless of task type. While such focused routing may encourage sharper specialization, it also limits expert diversity, wastes model capacity, and increases vulnerability to under-performing experts. Other methods SMoE, $\sigma$-MoE, MoE++, and SharedE-V2/V3 fall in an intermediate regime. They consistently show higher EWA for reasoning tasks compared to simpler ones, indicating that the router dynamically allocates weight more broadly when solving complex problems. This behavior suggests that harder tasks encourage the router to leverage complementary knowledge from multiple experts, while simpler tasks tend to concentrate on fewer experts.

### v3

@@ 3.3 Pre-training LLM: Decoder-only Switch Transformer
A key design of our language modeling experiment is to establish a standardized yet flexible setup that enables fair comparison across different MoE algorithms while maintaining computational efficiency. We adopt a decoder-only Transformer architecture in the style of Switch Transformer (Fedus et al., 2022), where the feedforward sublayer is replaced by sparsely activated MoE modules. This design supports scaling from small (0.15B) to large (0.68B) models and provides a unified backbone for evaluating MoE variants. To ensure consistency, we keep the architecture and optimization pipeline fixed across experiments, varying only the routing strategies.
Training pipeline. Our experimental setup builds upon the framework of Csordás et al. (2023), with adjustments tailored to our model configurations and minor refinements to the baseline hyperparameters. Our codebase supports all major forms of parallelism, including data, tensor, pipeline, and expert parallelism, ensuring scalability across model sizes. MoE variants are implemented by modularly replacing the MoE component, while keeping other training components identical. This design ensures consistency and reproducibility across experiments, allowing us to isolate the impact of routing. Full specifications of model scales, datasets, and hyperparameters are provided in Appendix A.2 and Appendix B.2.
@@ 3.5 Design Principles
LibMoE is designed with a strong emphasis on modularity, extensibility, and practical usability to accelerate research on SMoE algorithms. At its core, LibMoE consists of three major modules. The MoE module provides a unified abstraction for routing strategies, supporting customizable designs such as gating mechanisms, load balancing losses, and expert selection logic. The training module handles optimization workflows across both LLM and VLM tasks, allowing researchers to integrate pretrained backbones, apply sparse upcycling or pretraining from scratch, and utilize distributed strategies such as data parallelism and model sharding. Finally, the evaluation module enables seamless plug-and-play evaluation, supporting a wide range of benchmarks and task formats through standardized APIs.
@@ 4.1 SMoE Algorithms
To demonstrate that LibMoE is well-suited for SMoE research, we implement and benchmark seven state-of-the-art SMoE algorithms. First, we consider methods that modify or improve the router network. These include the original SMoE (Fedus et al., 2022), which uses a softmax router; the sigmoid-gated SMoE (Csordás et al., 2023), which has been shown to achieve faster convergence; and XMoE (Chi et al., 2022), which mitigates representation collapse and enhances routing capacity through input dimension reduction and cosine-normalized routing. Second, we incorporate recent designs that utilize shared expert mechanisms, such as DeepSeek-V2 (Liu et al., 2024a) with softmax routing and DeepSeek-V3 (Liu et al., 2024c) with a sigmoid router. Inspired by these designs, we implement SharedE-V2 and SharedE-V3, which adopt the shared-expert principles of DeepSeek-V2 and DeepSeek-V3, respectively. Lastly, we include MoE++(Jin et al., 2024), which augments standard SMoE with zero-computation experts and pathway-aware routing to improve efficiency and throughput, and TC-MoE(Yan et al., 2025), which introduces a ternary expert choice space and reward-based routing to reduce redundancy while enhancing model expressiveness. These diverse algorithms allow us to evaluate LibMoE’s compatibility with a wide range of routing strategies and architectural innovations.
@@
MoE Method AI2D Text VQA GQA MM Bench Hallusion Bench Math Vista MMMU MMStar Pope MME MME RW OCR Bench AVG Acc $\uparrow$ AVG Rank $\downarrow$ LLAVA + OneVision / 1M2 samples SMoE 69.56 43.93 61.51 71.31 46.90 37.90 41.56 41.23 86.28 63.33 27.83 37.50 52.40 5.96 XMoE 69.72 43.93 61.52 72.25 47.42 38.50 42.11 43.99 86.61 63.81 29.18 39.40 53.20 3.33 $\sigma$-MoE 69.79 44.69 61.70 71.74 47.00 38.40 43.11 42.08 86.69 63.80 29.70 38.40 53.09 3.33 SharedE-V2 70.56 45.04 61.34 71.13 47.11 39.50 42.78 42.73 86.53 63.93 29.55 38.70 53.24 3.33 SharedE-V3 71.92 44.59 61.93 72.59 46.37 38.20 42.33 43.30 86.78 64.61 29.29 39.40 53.44 2.46 TC-MOE 70.08 43.75 61.89 71.05 45.74 38.10 41.89 43.64 86.76 63.05 31.84 38.30 53.01 4.50 MOE++ 70.13 43.37 61.52 71.39 46.16 38.60 40.78 43.24 86.60 63.26 28.19 37.50 52.56 5.08 LLAVA / 665K samples SMoE 65.52 41.51 61.62 72.25 41.75 29.50 41.67 42.24 87.13 61.05 32.15 31.70 50.67 3.42 XMoE 65.84 41.96 61.61 72.16 41.85 31.60 41.67 40.32 86.64 60.95 32.73 33.20 50.88 3.08 $\sigma$-MoE 65.52 42.08 61.74 71.22 40.80 30.30 41.22 41.99 86.64 61.44 32.31 32.50 50.65 3.71 SharedE-V2 64.77 41.96 61.27 71.74 41.85 30.90 43.33 41.56 86.82 60.52 32.88 31.40 50.75 3.67 SharedE-V3 65.58 42.06 61.26 72.42 41.43 30.60 42.44 41.75 86.81 60.93 31.47 32.60 50.86 3.33 TC-MOE 65.50 40.70 61.19 71.22 42.06 29.30 41.22 41.10 86.53 60.30 31.89 33.00 50.34 5.42 MOE++ 65.03 41.44 60.61 71.74 42.69 30.30 43.00 40.32 86.58 60.11 31.32 31.20 50.36 5.38
Table 2: Performance of SMoE variants with 66 total experts and 8 active experts in the language pre-training setting, evaluated on small-scale (0.15B) and large-scale (0.68B) models. PPL denotes perplexity for language modeling. Bold values denote the best result in each column; $\downarrow$$\uparrow$ indicate that lower/higher values are preferable, respectively.
@@
MoE Method PPL $\downarrow$ LAM BADA BLiMP CBT Hella Swag PIQA ARC- Easy RACE SIQA Common SenseQA AVG Acc $\uparrow$ AVG Rank $\downarrow$
SMoE 13.63 25.27 77.71 84.18 29.43 57.94 32.68 30.11 35.62 24.65 44.18 4.50
XMoE 13.98 24.57 76.53 84.12 29.34 58.27 32.26 29.69 35.47 24.49 43.86 6.45
$\sigma$-MoE 13.61 25.43 77.38 84.23 29.13 58.92 32.73 31.05 34.90 24.90 44.30 4.30
SharedE-V2 13.49 25.29 77.37 84.33 29.38 60.17 33.83 31.02 35.57 24.98 44.66 2.80
SharedE-V3 13.42 25.49 77.20 84.40 29.38 59.14 32.52 30.60 35.57 25.47 44.42 3.20
TC-MoE 13.51 25.60 76.91 84.68 29.27 59.03 33.02 30.63 36.03 26.37 44.62 2.90
Small Model (0.15B) MoE++ 13.54 25.45 77.23 84.83 29.28 58.49 33.49 30.11 35.62 24.49 44.33 3.85
SMoE 9.51 37.13 80.47 89.83 37.49 64.36 38.22 33.03 37.41 26.54 49.39 5.15
XMoE 9.66 35.25 80.38 89.35 37.19 64.20 38.99 32.95 37.77 28.34 49.38 5.75
$\sigma$-MoE 9.46 37.56 81.08 89.57 37.52 64.91 39.15 32.68 37.67 28.50 49.85 3.60
SharedE-V2 9.52 37.11 80.98 89.93 37.14 64.36 38.06 33.17 36.95 27.35 49.45 5.25
SharedE-V3 9.49 36.88 81.28 89.65 37.32 65.72 38.86 33.12 38.59 28.09 49.95 3.60
TC-MoE 9.38 37.87 81.21 90.19 37.95 64.47 39.28 33.77 37.92 27.85 50.06 2.35
Large Model (0.68B) MoE++ 9.38 38.80 80.88 89.77 37.70 64.64 39.37 34.02 37.97 28.34 50.16 2.30
4.2 Performance Comparison
Across both vision–language modeling and language modeling settings, we observe a consistent trend: performance differences among contemporary SMoE variants remain relatively modest under matched training conditions. In the VLM sparse upcycling regime, no single method consistently outperforms others, with only marginal variations observed across benchmarks. Similarly, in language model pretraining, while recent MoE designs exhibit incremental improvements, these gains do not amount to a decisive advantage over the standard SMoE formulation. Collectively, these results reflect current practical deployments of frontier models, where the original SMoE design is often preferred for its simplicity, stability, and scalability, as opposed to more complex routing mechanisms that yield limited empirical benefits under comparable training conditions.
@@ 5 Analysis
Beyond empirical performance, we analyze how different MoE architectures route, specialize, and balance experts during training. We focus on expert selection dynamics, capacity utilization, and the effects of initialization, architecture, and task domain. Our analysis covers both from-scratch pretraining (0.15B LLM on 6.85B tokens) and sparse upcycling (VLM fine-tuned on LLaVA-665K), providing a unified view of routing behavior across regimes.
@@
This experiment investigates how the amount of training data affects expert selection dynamics across SMoE algorithms. We quantify routing stability using the Expert Change Rate (ECR). Concretely, for an evaluation set $\mathcal{D}$}, we run two checkpoints (at steps $T$ and T+1T{+}1) and extract the selected expert for every token at every MoE layer. We compute $\mathrm{ECR}_{T\rightarrow T{+}1}({\mathcal{D}})$ as the proportion of layer–token assignments whose selected expert differs between the two checkpoints. A low ECR indicates that routing has largely converged, consistently mapping the same inputs to the same experts, whereas a high ECR reflects unstable routing. Because $\mathcal{D}$} is held fixed across checkpoints, ECR can be interpreted as a form of routing regret: persistently high values late in training indicate that the router continues to revise assignments for the same inputs, while a rapid decay indicates early convergence to a stable partition of the input space and supports expert specialization.
@@
Figure 2 reports ECR throughout training for both vision-language models (MME, MathVista; sparse upcycling) and language models (Validation, HellaSwag; from-scratch pretraining). Across all settings, we observe that SMoE algorithms exhibit a consistently decreasing ECR as training progresses. Among the evaluated methods, XMoE shows stronger late-stage convergence on VLM tasks. We attribute this to its cosine-normalized routing and the smaller learning rate, which makes this setup more efficient. In contrast, during language-model pretraining, XMoE exhibits a relatively high ECR (peaking around the $40\%$ training mark), consistent with the degraded performance reported in Table 2. We hypothesize that this degradation arises from an interaction between cosine-normalized routing and the pretraining learning rate. By bounding the router logits (up to a temperature/scale), cosine/L2 normalization can reduce the effective logit margin and make softmax gating more sensitive to the chosen scale and step size (Agarwala et al., 2020). Moreover, since normalization choices can affect gradient scaling, an overly large learning rate can exacerbate optimization instability and even trigger divergence accompanied by growing internal activations/norms (Xiong et al., 2020; Rybakov et al., 2024).
Furthermore, we observe a divergent trend in the behavior of shared expert architectures namely, SharedE-V2 and SharedE-V3 across different tasks. In vision language models (VLMs), both variants exhibit significantly higher routing stability compared to other methods, likely because the router effectively selects among only KSK{-}S non-shared experts (with $S$ shared experts always active), which reduces the degrees of freedom in expert assignment. However, this advantage disappears during language model pretraining. We hypothesize that this divergence arises from the distinct initialization schemes and functional roles of shared experts. In the VLM sparse upcycling setup, SMoE layers are initialized from a pretrained dense model, where shared experts already encode general purpose visual linguistic representations. These pretrained experts act as stable anchors, enabling the router to focus its capacity on distributing tokens among non-shared experts, thereby reducing volatility in expert selection. In contrast, language model pretraining starts from random initialization. Here, both the router and shared experts must co-adapt from scratch. As the shared experts’ representations evolve over time, their influence on the final output changes accordingly. This compels the router to continuously revise its decisions not just for load balancing but also to adapt to the shifting behavior of shared experts in order to minimize the task loss. The result is a tight coupling between unstable shared experts and routing dynamics, which leads to higher expert change rates and delayed convergence.
@@
Taken together, these findings highlight that routing stability is shaped by the training regime. In from-scratch language-model pretraining, both the router and experts must co-adapt under random initialization, so early-stage stability is critical to prevent prolonged assignment churn. In contrast, VLM sparse upcycling starts from a pretrained dense checkpoint, where experts can serve as stable general-purpose anchors; consequently, routing tends to evolve more smoothly and robustness becomes more important than rapid early convergence. Therefore, effective sparse MoE designs should account for both regimes: fast stabilization during from-scratch pretraining and sustained robustness during upcycling.
@@
Figure 4: Performance of SMoE variants under routing perturbations, where the top-1 expert is replaced by the top-(K+1)(K{+}1) expert in vision–language and language modeling tasks.
To assess whether SMoE algorithms fully exploit their expert capacity, we conduct a diagnostic routing perturbation experiment across seven architectures. Specifically, we intentionally degrade the routing decision by replacing the top-1 expert (i.e., the expert with the highest affinity score) with the top-(K+1)(K{+}1) expert, and quantify the resulting performance change using $\Delta\text{Total Score}$ (see Appendix E.1 for details). Here, a higher $\Delta\text{Total Score}$ indicates improved performance relative to the original routing configuration. This controlled perturbation allows us to systematically measure each model’s sensitivity to suboptimal expert assignments and to identify potential inefficiencies in expert utilization. Figure 4 reports the cumulative accuracy drop across benchmarks under two settings: replacing only the top-1 expert (DropTop1) and replacing both the top-1 and top-2 experts (DropTop1&\&2). As expected, most SMoE variants exhibit performance degradation, confirming that precise expert selection is critical to maintaining model quality. Overall, in the language modeling (pretraining) setting, all methods suffer large and relatively uniform performance drops, reflecting a consistent reliance on precise expert routing across architectures. In contrast, vision-language models exhibit greater divergence. Notably, SMoE shows a counterintuitive a +2.3 gain in $\Delta\text{Total Score}$ under the DropTop1 perturbation suggesting that its original routing configuration may have failed to effectively utilize its expert capacity. This exposes a potential inefficiency in the router, where suboptimal expert choices actually yield better outcomes. Conversely, TC-MoE suffers the most severe degradation, highlighting its strong dependence on highly optimized, task-aligned routing decisions.
In summary, this experiment highlights that the impact of routing quality is fundamentally intertwined with architectural design across SMoE variants. Architectures such as TC-MoE and MoE++ exhibit pronounced performance degradation under routing perturbations, indicating that their gains rely heavily on precise and task-aligned expert selection. This sensitivity reveals a strong coupling between routing robustness and the effectiveness with which the architecture exploits its representational capacity. However, the results also demonstrate that routing robustness alone is insufficient to guarantee the highest absolute performance. Each SMoE variant is constrained by an inherent architectural capacity ceiling, beyond which improvements in routing precision yield diminishing returns. Consequently, optimal performance depends not only on selecting the “right” experts, but on whether the routing mechanism is capable of fully leveraging the architectural capacity available. These findings suggest that routing behavior should be interpreted in conjunction with architectural design, rather than as an isolated indicator of model quality.
@@
In Figure 5, we examine the behavior of the expert selection mechanism by measuring how frequently each expert is activated across subtasks in the MME benchmark, using the Expert Allocation Entropy (EAE) metric (see Appendix E.2). This analysis enables us to assess the degree of specialization exhibited by individual experts across different domains.
@@
Overall, these findings reveal a spectrum of expert allocation behaviors from highly specialized but potentially imbalanced (e.g., TC-MoE), to adaptive and task-aware (e.g., XMoE), to balanced yet under-specialized (e.g., SMoE family) highlighting the importance of routing strategies that balance specialization with generalization.
@@
In Figure 6, we report the normalized entropy of expert-weight allocation (EWA), calculated as the Shannon entropy of the softmax router weights over selected experts, and normalized with the maximum entropy of a vector with $K$ elements and averaged over samples within each task group OCR, Coarse-Grained, and Reasoning to reveal trends in how confidently or evenly each routing strategy distributes weight across experts, which EWA larger than weights more uniform. Notably, XMoE exhibits the most desirable profile across all three axes of evaluation. First, it achieves the highest EWA, indicating that the router distributes weight relatively evenly among the selected experts. Second, the entropy remains remarkably consistent across all task groups, with nearly identical scores for Coarse-Grained, Fine-Grained, OCR, and Reasoning. This reflects a routing strategy balance weights to experts reduce appear some expert nominate and each expert contribution to output is nearly equal. At the other extreme, TC-MoE records the lowest EWA across all tasks, with minimal inter-task variation. This indicates that the router consistently concentrates weight on a small subset of experts, regardless of task type. While such focused routing may encourage sharper specialization, it also limits expert diversity, wastes model capacity, and increases vulnerability to under-performing experts. Other methods, including SMoE, $\sigma$-MoE, MoE++, and SharedE-V2/V3, fall into an intermediate regime. They consistently exhibit higher EWA on reasoning tasks than on simpler tasks, indicating that the router distributes weights more broadly across experts when handling complex problems. In contrast, for simpler tasks, routing decisions tend to concentrate weight on a smaller subset of experts.
@@
Router margin measures the dominance of the top-ranked expert in a Mixture-of-Experts (MoE) routing decision. It is defined as the difference between the top-1 and top-2 routing scores, with a larger margin indicating a stronger preference for the top-1 expert (see Appendix E.3). Figure 7 shows the layer-wise router margin across all MoE methods, revealing a consistent trend of increasing margin with model depth in both language modeling and vision–language tasks. This indicates that expert dominance becomes more pronounced in deeper layers. In particular, the router margin is close to zero in layer 0 for most methods, reflecting early training stages where routing is uncertain and experts receive balanced weights. As depth increases, the rising margin signals a transition from generalist processing toward more specialized expert allocation, with the exception of the final layer in vision–language models, where the pattern deviates.